# AlphaViT: A Flexible Game-Playing AI for Multiple Games and Variable Board Sizes

### v2

@@ 6 Conclusion and discussion
1 AlphaZero
AlphaZero integrates a deep neural network (DNN) with Monte Carlo Tree Search (MCTS), as illustrated in Fig. 1. The DNN processes input representing the current board state and the player, producing an estimated state value and move probabilities. MCTS then uses these outputs to identify the optimal move. Similarly, AlphaViT, AlphaViD, and AlphaVDA adopt this foundational framework.
1.1 Deep neural network in AlphaZero
AlphaZero’s deep neural network (DNN) predicts the value $v(s)$ and the move probability $\bm{p}$ with components $p(a|s)$ for each action $a$, given a state $s$. In the board game context, $s$ and $a$ represent the board state and the move, respectively. In the context of board games, $s$ represents the board state, and $a$ denotes a move. The DNN receives input representing the current board state and the current player’s disc color. Fig. A1 illustrates the DNN architecture, which consists of a Body (residual blocks) and a Head (value and policy heads). The value head outputs the estimated state value $v(s)$, while the policy head produces the move probabilities $\bm{p}$.
The input to the DNN is an $H\times W\times(2T+1)$ image stack that contains $12T+$ 1 binary feature planes of size $H\times W$. Here, $H\times W$ refers to the board size, and $T$ is the number of histories. The first $T$ feature planes represent the occupancy of the player’s discs, with a feature value of 1 indicating that a disc occupies the corresponding cell, and 0 otherwise. Similarly, the following $T$ feature planes represent the occupancy of the opponent’s discs. The last feature plane represents the disc color of the current player, with the disc colors of the first and second players represented by 1 and -1, respectively.
Figure A1: Deep neural network of AlphaZero.
1.2 Monte Carlo tree search in AlphaZero
@@
In the Select, the tree is searched from the root node $s_{\mathrm{root}}$ to the leaf node $s_{L}$ at time step $L$ using a variant of the PUCT algorithm. At each time step $t<L$, the selected action $a_{t}$ has the maximum score, as described by the following equation:
$a_{t}=\mathop{\rm arg~{}max}\limits_{a}(Q(s_{t},a)+C_{\mathrm{puct}}P(s_{t},a)\frac{\sqrt{N(s_{t})}}{1+N(s_{t},a)}),$ (9)
where $N(s_{t})$ is the number of parent visits and $C_{\mathrm{puct}}$ is the exploration rate. In this study, $C_{\mathrm{puct}}$ is constant, whereas in the original AlphaZero, $C_{\mathrm{puct}}$ increases slowly with search time. Additionally, when the parent node is the root node, the node selection is performed using an $\epsilon$-greedy algorithm based on the UCT scores.
In the Expand and Evaluate, the DNN evaluates the leaf node and outputs $v(s_{l})$ and $\bm{p}_{a}(s_{l})$. If the leaf node is a terminal node, $v(s_{L})$ is the color of the winning player’s disc. The leaf node is expanded and each edge $(s_{L},a)$ is initialized to $\{N(s_{L},a)=0,W(s_{L},a)=0,Q(s_{L},a)=0,P(s_{L},a)=p_{a}\}$.
In the Backup, the visit counts and values are updated in a backward pass through each step, $t\leq L$. The visit count is incremented by 1, $N(s_{t},a_{t})\leftarrow N(s_{t},a_{t})+1$, and the cumulative and average values are updated, $W(s_{t},a_{t})\leftarrow W(s_{t},a_{t})+v$, $Q(s_{t},a)\leftarrow W(s_{t},a_{t})/N(s_{t},a_{t})$.
Finally, in the Play, AlphaZero selects the action corresponding to the most visited edge from the root node.
2 Training
AlphaViT, AlphaViD, AlphaVDA, and AlphaZero use the same training scheme. The training process consists of three main components: Self-play, Augmentation, and Update, which are iterated $N_{\mathrm{iter}}$ times. This training algorithm is a modified version of the original AlphaZero, adapted to allow training on a single computer.
During the Self-play phase, an agent plays against itself $N_{\mathrm{self}}$ times. For the first $T_{\mathrm{opening}}$ turns, actions are stochastically selected from valid moves based on the softmax policy as follows:
$p(a\mid s)=\exp(N(s,a)/\tau)/\Sigma_{b}\exp(N(s,b)/\tau),$ (10)
where $\tau$ is a temperature parameter that controls the exploration. This stochastic exploration enables the agent to explore new and potentially better actions. After $T_{\mathrm{opening}}$, the most visited action is selected. Through Self-play, we collect board states, winners, and search probabilities. The search probabilities represent the probabilities of selecting valid moves at the root node in MCTS.
@@ 2 Training
During the Update phase, the deep neural network (DNN) is trained using mini-batch stochastic gradient descent with $N_{\mathrm{batch}}$ batch size and $N_{\mathrm{epochs}}$ epochs. The optimization process includes momentum and weight decay. The loss function $l$ combines the mean squared error between the predicted value $v$ and the winner’s disc color $c_{\mathrm{win}}$, and the cross-entropy loss between the search probabilities $\bm{\pi}$ and the predicted move probabilities $\bm{p}$:
@@
To train multiple games simultaneously, mini-batches are generated from the respective training data queue of each game. During the Update phase, mini-batches are sampled from these individual queues and used to update the DNN. For example, when an agent simultaneously trains Connect4, Gomoku, and Othello, a mini-batch of Connect4, a mini-batch of Gomoku, and a mini-batch of Othello are used orderly for the update.
3 Parameters
The hyperparameters for AlphaViT, AlphaViD, AlphaVDA, and AlphaZero are detailed in Table A1. The weight decay and momentum values are consistent with those specified in the AlphaZero pseudocode (Silver et al., 2018). All other parameters for AlphaZero are consistent with the previous implementation (Fujita, 2022). The hyperparameters of the other models were carefully hand-tuned to optimize their performance.
Table A2 lists the game-specific hyperparameters for AlphaViT, ALphaViD, AlphaVDA, and AlphaZero. The number of MCTS simulations ($N_{\mathrm{sim}}$) ranges from 200 to 400, depending on the game and board size. The number of self-play games per iteration is set to 30 for Connect4 variants and 10 for Gomoku and Othello variants. The opening phase ($T_{\mathrm{opening}}$) specifies the number of initial moves using softmax decision-making with a temperature parameter ($\tau$) that is adjusted based on the game and board size. For AlphaViT, AlphaViD, and AlphaVDA, we employ the AdamW optimizer in PyTorch, with all parameters set to their default values except for the learning rate. These hyperparameters were also carefully hand-tuned.
@@ 3 Parameters
Optimizer AdamW AdamW AdamW SGD
@@
Learning rate 0.0001 0.0001 0.0001 0.05
Momentum - - - 0.9
Weight decay 0.01 0.01 0.01 0.0001
Patch size 5 5 5 -
Stride for making patches 1 1 1 -
@@
Dim feedforward of encoder 1024 1024 1024 -
Num of encoder heads 8 8 8 -
Size of positional embeddings $512\times 256$ $512\times 256$ $512\times 256$
Num of decoder layers - 1 1 -
@@
Dim feedforward of decoder - 1024 1024 -
Num of decoder head - 8 8 -
@@
Connect4 Connect4 5x4 Gomoku Gomoku 66 Othello Othello 66
Num of simulations 200 200 400 200 400 200
Num of self-play 30 30 10 10 10 10
$T_{\mathrm{opening}}$ 4 4 8 4 8 4
@@
4 Monte Carlo tree search
MCTS is a best-first search method that does not require a predefined evaluation function (Winands, 2017). It operates through four strategic steps: Selection step, Playout step, Expansion step, and Backpropagation step (Winands, 2017). In the Selection step, the tree is traversed from the root node to a leaf node. Child node $c$ of parent node $p$ is selected to maximize the following score:
$\mathrm{UCT}=q_{c}/N_{c}+C\sqrt{\frac{\ln(N_{p}+1)}{N_{c}+\varepsilon}},$ (12)
where $q_{c}$ is the cumulative value of $c$, $N_{p}$ is the visit count of $p$, $N_{c}$ is the visit count of $c$, and $\varepsilon$ is a constant value to avoid division by zero. In the Playout step, a valid move is selected at random until the end of the game is reached. In the Expansion step, child nodes are added to the leaf node when the visit count of the leaf node reaches $N_{\mathrm{open}}$. In the Backpropagation step, the result of the playout is propagated along the path from the leaf node to the root node. If the MCTS player itself wins, loses, and draws, $q_{i}\leftarrow q_{i}+1$, $q_{i}\leftarrow q_{i}-1$, and $q_{i}\leftarrow q_{i}$, respectively. $q_{i}$ is the cumulative value of node $i$. The sequence of these four steps constitutes a single simulation. The simulation process is repeated $N_{\mathrm{sim}}$ times. After all simulations, MCTS selects the action corresponding to the most visited child node of the root. In this study, $C=0.5$, $\varepsilon=10^{-7}$, and $N_{\mathrm{open}}=5$.
5 Minimax algorithm
The minimax algorithm is a fundamental game-tree search technique that determines the optimal action by evaluating the best possible outcome for the current player. Each node in the tree contains a state, player, action, and value. The algorithm creates a game tree with a depth of $N_{\mathrm{depth}}=3$. The root node corresponds to the current state and minimax player. Next, the states corresponding to leaf nodes are evaluated. Then, the algorithm propagates the values from the leaf nodes to the root node. If the player corresponding to the node is the opponent, the value of the node is the minimum value of its child nodes. Otherwise, the value of the node is the maximum value of its child nodes. Finally, the algorithm selects the action corresponding to the root’s child node with the maximum value. The evaluation of leaf nodes is tailored to each game.
@@ 5 Minimax algorithm
For the Othello variants, the value of a node is calculated using the following equation:
@@
where $v_{(}x,y)$ is the value of cell $(x,y)$ and $o(x,y)$ is the occupancy of cell $(x,y)$. For 6x6 Othello and 8x8 Othello, the minimax algorithm evaluates each cell using Eq. 14 and 15, respectively. $o(x,y)$ is $1$, $-1$, and 00 if cell $(x,y)$ is occupied by the first player’s disc, the second player’s disc, and empty, respectively. For the Othello variants, the algorithm expands the tree to the terminal nodes after the last six turns. The terminal nodes have a value of $E_{\mathrm{end}}=1000c_{\mathrm{win}}c_{\mathrm{minimax}}$.
@@ 6 Elo rating
Elo rating is a widely used metric for evaluating the relative performance of players in two-player games. This allows us to estimate the probability of one player defeating another based on their current ratings. Given two players A and B with Elo ratings $e(A)$ and $e(B)$, respectively, the probability that player A will defeat player B, denoted $p(A\mathrm{defeats}B)$, is calculated using the following formula:
$p(A~{}\mathrm{defeats}~{}B)=1/(1+10^{(e(B)-e(A))/400}).$ (16)
@@
After a series of $N_{G}$ games between players A and B, player A’s Elo rating is updated to a new value $e^{\prime}(A)$ based on their performance:
$e^{\prime}(A)=e(A)+K(N_{\mathrm{win}}-N_{G}\times p(A~{}\mathrm{defeats}~{}B)),$ (17)
@@
where $N_{\mathrm{win}}$ is the number of times player A has won, and $K$ is a factor that determines the maximum rating adjustment after a single game. In this study, $K=8$.

### v3

@@ Acknowledgments
A AlphaZero
AlphaZero integrates a deep neural network (DNN) with Monte Carlo Tree Search (MCTS), as illustrated in Figure 1. The DNN processes input representing the current board state and the current player, producing an estimated state value and a move-probability vector. MCTS then uses these outputs to select the optimal move. This same framework is adopted by AlphaViT, AlphaViD, and AlphaVDA.
A.1 Deep Neural Network in AlphaZero
AlphaZero’s DNN predicts a value $v(s)$ and a move-probability vector $\bm{p}(s)$ with components $p(a\mid s)$ for each action $a$, given a state $s$. In the board game context, $s$ and $a$ represent the board state and the move, respectively. The DNN receives input representing the current board state and the current player’s disc color. Figure A1 illustrates the DNN architecture, which consists of a Body (residual blocks) and two Heads (value and policy heads). The value head outputs the estimated state value $v(s)$, while the policy head produces the move probabilities $\bm{p}(s)$.
[…]
Figure A1: DNN of AlphaZero.
A.2 Monte Carlo Tree Search in AlphaZero
@@
In Select, the tree is searched from the root node $s_{\mathrm{root}}$ to the leaf node $s_{L}$ at time step $L$ using a variant of the PUCT algorithm. At each time step $t<L$, the selected action $a_{t}$ has the maximum score, as described in Eq. 9:
$a_{t}=\mathop{\rm arg\leavevmode\nobreak\ max}\limits_{a}(Q(s_{t},a)+C_{\mathrm{puct}}p(s_{t},a)\frac{\sqrt{N(s_{t})}}{1+N(s_{t},a)}),$ (9)
where $N(s_{t})$ is the number of parent visits and $C_{\mathrm{puct}}$ is the exploration rate. In this study, $C_{\mathrm{puct}}$ is constant, whereas in the original AlphaZero, $C_{\mathrm{puct}}$ increases slowly with search time Silver et al., (2018). Additionally, when the parent node is the root node, the node selection is performed using an $\epsilon$-greedy algorithm based on the UCT scores.
In Expand and Evaluate, the DNN evaluates the leaf node and outputs $v(s_{L})$ and $\bm{p}(s_{L})$. If the leaf node is a terminal node, $v(s_{L})$ is the color of the winning player’s disc. The leaf node is expanded and each edge $(s_{L},a)$ is initialized to $\{N(s_{L},a)=0,W(s_{L},a)=0,Q(s_{L},a)=0,p(s_{L},a)=p(a\mid s_{L})\}$.
In Backup, the visit counts and values are updated for each step $t\leq L$ during the backward pass. The visit count is incremented by 1, $N(s_{t},a_{t})\leftarrow N(s_{t},a_{t})+1$, and the cumulative and average values are updated, $W(s_{t},a_{t})\leftarrow W(s_{t},a_{t})+v$, $Q(s_{t},a_{t})\leftarrow W(s_{t},a_{t})/N(s_{t},a_{t})$.
Finally, in Play, AlphaZero selects the action corresponding to the most visited edge from the root node.
B Training
AlphaViT, AlphaViD, AlphaVDA, and AlphaZero share the same three-stage training loop: Self-play, Augmentation, and Update. One complete cycle of the three stages is called an iteration and is repeated $N_{\mathrm{iter}}$ times. This training algorithm is a modified version of the original AlphaZero, adapted for a single-machine setting.
During the Self-play phase, an agent plays against itself $N_{\mathrm{self}}$ times. For the first $T_{\mathrm{opening}}$ turns, actions are stochastically selected among the valid moves according to the softmax policy defined in Eq. 10:
$p(a\mid s)=\exp(N(s,a)/\tau)/\sum_{b}\exp(N(s,b)/\tau),$ (10)
where $\tau$ is a temperature parameter that controls the exploration. This stochastic exploration enables the agent to explore new and potentially better actions. After $T_{\mathrm{opening}}$, the most visited action is selected. During Self-play, we record the board states, game outcomes (winners), and the search probabilities. The search probabilities represent the probabilities of selecting valid moves at the root node in MCTS.
@@ B Training
During the Update phase, the DNN is trained using mini-batch stochastic gradient descent with a batch size of $N_{\mathrm{batch}}$ for $N_{\mathrm{epochs}}$ epochs. The optimizer uses weight decay. The loss function $l$ combines the mean squared error between the predicted value $v$ and the winner’s disc color $c_{\mathrm{win}}$, and the cross-entropy loss between the search probabilities $\bm{\pi}$ and the predicted move probabilities $\bm{p}$. The loss function is defined in Eq. 11:
@@
To train multiple games simultaneously, mini-batches are generated from the respective training data queue of each game. During the Update phase, mini-batches are sampled from these individual queues and used to update the DNN. For example, when an agent simultaneously trains Connect 4, Gomoku, and Othello, we sequentially use one mini-batch from the training data queues of Connect 4, Gomoku, and Othello to update the network.
C Parameters
The hyperparameters for AlphaViT, AlphaViD, AlphaVDA, and AlphaZero are listed in Table A1. For training, we employ the AdamW optimizer in PyTorch, with all parameters set to their default values except for the learning rate. All other parameters for AlphaZero were consistent with the previous implementation (Fujita,, 2022). The hyperparameters of the other models were carefully hand-tuned to optimize their performance.
Table A2 lists the game-specific hyperparameters for AlphaViT, AlphaViD, AlphaVDA, and AlphaZero. The number of MCTS simulations ($N_{\mathrm{sim}}$) ranges from 200 to 400, depending on the game and board size. The number of self-play games per iteration is set to 30 for Connect 4 variants and 10 for Gomoku and Othello variants. The opening phase ($T_{\mathrm{opening}}$) specifies the number of initial moves using softmax decision-making with a temperature parameter ($\tau$) that is adjusted based on the game and board size.
@@ C Parameters
Optimizer AdamW AdamW AdamW AdamW
@@
Learning rate 0.0001 0.0001 0.0001 0.0001
Weight decay 0.01 0.01 0.01 0.01
Patch size $\times 5$ 5 $\times 5$ 5 $\times 5$ 5 -
Patch stride 1 1 1 -
@@
Encoder feedforward dimension 1024 1024 1024 -
Number of encoder heads 8 8 8 -
Size of positional embeddings 512×256512$\times$ 256512 × 256 512×256512$\times$ 256512 × 256 512×256512$\times$ 256512 × 256 -
Number of decoder layers - 1 1 -
@@
Decoder feedforward dimension - 1024 1024 -
Number of decoder heads - 8 8 -
@@
Connect 4 Connect 4 5x4 Gomoku Gomoku 6x6 Othello Othello 6x6
Number of simulations 200 200 400 200 400 200
Number of self-play 30 30 10 10 10 10
$T_{\mathrm{opening}}$ 4 4 8 6 8 6
@@
D Monte Carlo tree search
MCTS is a sampling-based tree-search algorithm that does not require a predefined evaluation function (Browne et al.,, 2012; Winands,, 2017). It operates through four strategic steps: Selection step, Expansion step, Playout step, and Backpropagation step (Winands,, 2017). In the Selection step, the tree is traversed from the root node to a leaf node. Child node $c$ of parent node $p$ is selected to maximize the score defined in Eq. 12:
$\mathrm{UCT}=q_{c}/N_{c}+C\sqrt{\frac{2\ln(N_{p}+1)}{N_{c}+\varepsilon}},$ (12)
where $q_{c}$ is the cumulative value of $c$, $N_{p}$ is the visit count of $p$, $N_{c}$ is the visit count of $c$, and $\varepsilon$ is a constant value to avoid division by zero. In the Expansion step, child nodes are added to the leaf node when the visit count of the leaf node reaches $N_{\mathrm{open}}$. In the Playout step, a valid move is selected at random until the end of the game is reached. In the Backpropagation step, the result of the playout is propagated along the path from the leaf node to the root node. If the MCTS player itself wins, loses, and draws, $q_{i}\leftarrow q_{i}+1$, $q_{i}\leftarrow q_{i}-1$, and $q_{i}\leftarrow q_{i}$, respectively. $q_{i}$ is the cumulative value of node $i$. The sequence of these four steps constitutes a single simulation. The simulation process is repeated $N_{\mathrm{sim}}$ times. After all simulations, MCTS selects the action corresponding to the most visited child node of the root. In this study, $C=0.5$, $\varepsilon=10^{-7}$, and $N_{\mathrm{open}}=5$.
E Minimax algorithm
The minimax algorithm is a fundamental game-tree search technique that determines the optimal action by evaluating the best possible outcome for the current player. Each node in the tree contains a state, player, action, and value. The algorithm creates a game tree with a depth of $N_{\mathrm{depth}}=3$. For the Othello variants, the algorithm expands the tree to the terminal nodes after the last six turns. The root node corresponds to the current state and minimax player. Next, the states corresponding to leaf nodes are evaluated. Then, the algorithm propagates the values from the leaf nodes to the root node. If the player corresponding to the node is the opponent, the value of the node is the minimum value of its child nodes. Otherwise, the value of the node is the maximum value of its child nodes. Finally, the algorithm selects the action corresponding to the root’s child node with the maximum value. The evaluation of leaf nodes is tailored to each game.
@@ E Minimax algorithm
For the Othello variants, the value of a node is calculated using Eq. 13.
@@
[…]
@@ F.1 Calculation of Elo rating
Elo rating is a widely used metric for evaluating the relative performance of players in two-player games. This metric allows us to estimate the probability of one player defeating another based on their current ratings. Given two players A and B with Elo ratings $e(A)$ and $e(B)$, respectively, the probability that player A will defeat player B, denoted $p(A\mathrm{defeats}B)$, is calculated using Eq. 16:
$p(A\leavevmode\nobreak\ \mathrm{defeats}\leavevmode\nobreak\ B)=1/(1+10^{(e(B)-e(A))/400}).$ (16)
@@
After a series of $N_{G}$ games between players A and B, player A’s Elo rating is updated to a new value $e^{\prime}(A)$ based on their performance. The update is performed using Eq. 17:
$e^{\prime}(A)=e(A)+K(N_{\mathrm{win}}-N_{G}\times p(A\leavevmode\nobreak\ \mathrm{defeats}\leavevmode\nobreak\ B)),$ (17)
@@
where $N_{\mathrm{win}}$ denotes the total score accumulated by player A across the $N_{G}$ games (counting each win as 1, each draw as 0.5, and each loss as 0), and $K$ is a factor that determines the maximum rating adjustment after a single game. In this study, $K=8$.
F.2 Calculation of 95% confidence intervals for Elo ratings
To quantify estimation uncertainty, we compute 95% confidence intervals via percentile bootstrap (LMSYS Org,, 2023; Tang et al.,, 2025):
  1. (1)
    Resample the full set of $T$ game results with replacement from the original dataset (which contains $T$ samples), yielding a new dataset of size $T$.
  2. (2)
    Apply the same sequential-update Elo rule (initial rating = 1500, $K=8$) to compute ratings for all players.
  3. (3)
    Repeat steps 1 and 2 for $B=1000$ replicates to obtain a distribution of Elo ratings.
  4. (4)
    Define the 95% confidence interval for each player as the 2.5th and 97.5th percentiles of its bootstrap Elo distribution.
This yields point estimates from the original data and interval estimates from the bootstrap replicates, enhancing the statistical rigor of the evaluation.
F.3 Calculation of 95% confidence intervals for ratio of Elo ratings between deeper and shallower models
We compute an Elo rating for each training iteration ($t$) by replaying the complete match log in chronological order. For player $p$, we obtain an iteration series ${E_{p,t}}(t=$ 1,2, $\ldots$,10,20,40, $\ldots$,100,200,300, $\ldots$,3000))). To summarize playing strength over the final one thousand iterations, we consider the ten checkpoints in the interval $T=\{2100,2200,\dots,3000\}$. The mean Elo in this window is $\bar{E}_{p,s}=\frac{1}{|T|}\sum_{t\in T}E_{p,t}$, where ($s\in\{\text{shallower},\text{deeper}\}$) labels the shallower and deeper variants of the same architecture. From these averages we form the relative performance ratio $R=\bar{E}_{p,\text{deep}}/\bar{E}_{p,\text{shallow}}$. Additionally, we also compute the difference in average Elo ratings $\Delta=\bar{E}_{p,\text{deep}}-\bar{E}_{p,\text{shallow}}$. Sampling uncertainty for $R$ is quantified using a non-parametric percentile bootstrap: the ten checkpoint rows are resampled $B=1000$ times with replacement, the bootstrap statistics ($\{R^{*b}\}_{b=1}^{B}$) are recomputed, and a 95% confidence interval for $R$ is derived from the 2.5th and 97.5th percentiles of the resulting empirical distribution.
G Additional AlphaZero experiments
These experiments systematically evaluated how architectural depth and optimization methods influenced the learning dynamics and performance of AlphaZero models. The baseline configuration for this study, denoted as Res6, utilized a neural network architecture comprising six residual blocks trained with the AdamW optimizer. The deeper variant, Res10, increased the architectural depth to ten residual blocks and similarly employed AdamW optimization. A further variant, Res10 fine-tuning, represented the Res10 model refined through fine-tuning by employing the same procedure described in Subsection 5.4. Notably, before fine-tuning, only the heads were initialized with random weights because these heads for the small board configuration were not utilized and were replaced with new heads suitable for large board configurations. Finally, to specifically investigate the impact of optimizer selection, the Res6 SGD configuration maintained the six-block structure of Res6, but substituted AdamW with stochastic gradient descent (SGD). The SGD hyperparameters (the learning rate, weight decay, and momentum) were set to 0.05, 0.0001, and 0.9, respectively. The training process for Res6 SGD was identical to that denoted in Appendix B, except for the optimizer choice. Elo ratings were calculated according to the methodology detailed in Subsection 5.2.
Figure A2 illustrates the progression of Elo ratings over training iterations for the AlphaZero configurations across three games: Connect 4, Gomoku, and Othello, in large board scenarios. The results indicate that all AlphaZero variants achieve high Elo ratings with sufficient training. Notably, the Res10 fine-tuning configuration exhibits a rapid initial improvement in Gomoku, quickly attaining a high Elo rating relative to the non-fine-tuned Res10 configuration. Conversely, the Res6 SGD configuration consistently demonstrates inferior performance compared to its AdamW-optimized counterpart, plateauing prematurely and failing to achieve comparable Elo ratings across all games.
Collectively, these findings underscore the efficacy of Res6 architecture optimized with AdamW, affirming its suitability as a stable baseline for comparative studies. Moreover, AlphaZero demonstrates notable efficiency in terms of parameter count and computational resource requirements relative to transformer-based models, such as AlphaViT, AlphaViD, and AlphaVDA. In addition, deeper AlphaZero architectures benefit significantly from fine-tuning, rapidly achieving strong performance with reduced computational overhead.
Figure A2: Elo rating progression over training iterations for AlphaZero variants on large board configurations in Connect 4, Gomoku, and Othello. Res6: AlphaZero with six residual blocks. Res10: AlphaZero with ten residual blocks. Res10 fine-tuning: Res10 with fine-tuning. Res6 SGD: Res6 trained using the SGD optimizer.
H Additional MCTS experiment
We conducted an additional experiment to investigate the impact of varying simulation counts on the performance of MCTS in Connect 4, using the large board configuration. The MCTS algorithm was systematically evaluated using a fixed number of simulations ranging from 10 to 6400. Each MCTS variant was evaluated against the set of agents previously described (see Table 3), with the opponents’ Elo ratings maintained at the values already established. Elo ratings for the MCTS variants were calculated using the method described in Subsection 5.2.
The results of these evaluations are presented in Table A3, indicating a significant improvement in Elo ratings as the simulation count increases from 10 to 1600. Beyond 1600, the incremental Elo gain becomes less pronounced; 6400 adds only 30$\approx$ 30≈ 30 Elo, and 3200 shows no improvement, illustrating diminishing returns. This observed trend aligns closely with findings reported by Baier and Winands, (2015), who demonstrated that enhancing the inference time, conceptualized as the number of simulations in MCTS, does not necessarily lead to continuous performance improvements in Connect 4. Furthermore, we observe a substantial difference in performance between 100- and 400-simulation configurations, which are selected as baseline configurations.
Table A3: Elo ratings of MCTS variants in Connect 4
Simulations Elo rating 95% CI
10 1048.3 [1004.0, 1093.1]
100 1316.8 [1273.3, 1366.4]
200 1431.4 [1376.5, 1472.3]
400 1509.8 [1476.3, 1564.8]
800 1573.6 [1522.8, 1623.3]
1600 1676.7 [1611.0, 1715.0]
3200 1673.5 [1624.3, 1730.7]
6400 1708.4 [1674.1, 1767.1]
I Computational performance
Tables A4 and A5 summarize the computational performance of each model for Othello, including the number of parameters, peak GPU memory consumption, and mean execution time during training and inference, respectively. All experiments were conducted on a custom-made computer equipped with an Intel Core i9-12900K CPU, 64 GiB of DDR5 RAM, and two NVIDIA RTX 4060 Ti 16 GiB GPUs. For training, we employed a data-parallel approach utilizing both GPUs and ten parallel self-play games (five per GPU). During inference, a single process on a single GPU was used. To measure the training and inference speeds, two precision modes were considered: automatic mixed precision (FP16) and full precision (FP32). The experimental setup included PyTorch 2.3.1, NVIDIA Driver Version 560.35.03, and CUDA Version 12.6.
Peak GPU memory usage (in MiB) denotes the highest memory allocation recorded during training and inference. During training, peak memory usage represents the combined memory usage across both GPUs. Training speed was reported as seconds per iteration (s/iteration), where an iteration comprised one complete cycle of self-play, data augmentation, and network updates (as detailed in Appendix B). Inference speed was measured in seconds per decision (s/decision), reflecting the time required to compute one move. Mean values for training and inference speeds were computed over ten training iterations and ten full games played by the agent against itself, respectively.
The results demonstrate that for training, FP16 generally provides faster performance and reduced memory usage compared to FP32 for AlphaViT, AlphaViD, and AlphaVDA, but not for AlphaZero. Conversely, during inference, FP32 is generally faster and more memory-efficient than FP16 for all models. Consequently, FP16 was used for training and FP32 was used for inference in this study.
Additionally, we employed the number of model parameters and other hyperparameters (e.g., number of games per iteration and batch size) listed in Table A2 for training and evaluating the agents because our computational resources were limited to a single computer with two GPUs, and we were able to train and evaluate the agents with practical computation time for this research.
It is important to note that computation time is significantly influenced not only by model size but also by the game engine implementation, hardware specifications, and hyperparameters, such as the number of parallel processes for self-play and the number of simulations. Achieving a practical balance between these factors is crucial to ensure efficient training and inference.
Initially, for each proposed architecture, we selected the shallowest encoder depth (4 encoder layers for AlphaViT and one encoder layer for AlphaViD/AlphaVDA) that resulted in approximately 11 million parameters. This depth was chosen to achieve a balanced trade-off between model size, computational efficiency, and expected performance. However, for large board configurations, initial results indicated that these shallower models (especially AlphaViD and AlphaVDA) did not exhibit competitive performance against the AlphaZero baseline. Consequently, we introduced deeper models—AlphaViT with eight encoder layers (AlphaViT L8) and AlphaViD/AlphaVDA with five encoder layers (AlphaViD L5, AlphaVDA L5)—resulting in approximately 20 million parameters. These deeper configurations represented practical upper limits given our available computational resources, in particular the memory and performance constraints of our hardware setup consisting of two NVIDIA RTX 4060 Ti GPUs, each with 16 GiB of GPU memory.
Table A4: Computational performance during training
Model Params (M) Peak Mem (MiB) Mean time (sec/iteration)
FP16 FP32 FP16 FP32
AlphaViT L4 11.2 6672 9582 131.2 132.5
AlphaViT L8 19.6 11260 17292 197.3 208.8
AlphaViD L1 11.5 4980 7866 113.7 114.1
AlphaViD L5 19.9 9390 14902 180.6 189.1
AlphaVDA L1 11.3 4540 7074 109.6 110.6
AlphaVDA L5 19.8 9232 14598 177.8 184.9
AlphaZero (6 ResBlocks) 7.1 3526 3107 112.2 102.9
AlphaZero (10 ResBlocks) 11.8 5262 4448 149.4 134.3
 Peak memory is the maximum aggregated GPU usage observed over ten iterations. Mean iteration time (second/iteration) is averaged over the same interval.
Table A5: Computational performance during inference
Model Params (M) Peak Mem (MiB) Mean time (sec/decision)
FP16 FP32 FP16 FP32
AlphaViT L4 11.2 253 237 0.7541 0.6546
AlphaViT L8 19.6 307 261 1.038 0.6763
AlphaViD L1 11.5 251 239 0.7104 0.5802
AlphaViD L5 19.9 307 263 1.006 0.8018
AlphaVDA L1 11.3 249 237 0.7116 0.5943
AlphaVDA L5 19.8 305 261 0.9689 0.8199
AlphaZero (6 ResBlocks) 7.1 233 207 0.7864 0.7245
AlphaZero (10 ResBlocks) 11.8 273 227 0.9404 0.8293
Peak memory is the maximum single-GPU usage over ten self-play games. Mean decision time is averaged over the same games.